# Can Large Language Models Replace Data Scientists in Clinical Research?

### v1

@@ 2.4 Human-AI collaboration boosts productivity for data science in clinical research
The interface supports real-time interactions, allowing users to generate and execute code in a sandbox environment with instant visualizations. This removes the need for users to handle complex prompt crafting or manually switch between chat sessions and coding platforms like Jupyter Notebook. By simplifying the data science workflow, the platform empowers users with minimal coding expertise to perform complex data science tasks with ease.
@@
In our user study, we involved five medical doctors with varying levels of coding expertise. Each participant was assigned three studies [33, 34, 35], with approximately 10 coding tasks per study (Fig. 4d). Users worked with LLMs on our platform to complete these tasks and submitted their solutions once their code passed all the test cases. The difficulty levels of the tasks were quantified, with the distribution shown in Fig. 4b. During the study, we tracked two core actions: code generation and code improvement (debugging) requests. The statistics of these user behaviors are depicted in Fig. 4b, where most users completed the first two studies, and a few tackled the third. After the study, we analyzed the logs to compare the LLM-generated code with the final code solutions submitted by the users (Fig. 4a). Additionally, we conducted a survey to gather their feedback on the platform and their experience working with LLMs. The survey questions were built based on the Health Information Technology Usability Evaluation Scale (Health-ITUES) [36].
@@
The results of the code comparison analysis are presented in Fig. 4c, showing the distribution of the proportion of user-submitted code derived from LLM-generated solutions. We found that a significant portion of the user-submitted code was drawn from AI-generated code. For Easy tasks, the median proportions were 0.88, 0.87, and 0.84 across the three studies, indicating that users heavily relied on LLM-provided solutions when crafting their final submissions. For Medium and Hard tasks, the ratios were generally lower: in Study 1, the proportions were 0.44 for Medium tasks and 0.96 for Hard tasks, while in Study 2, the proportions were 0.75 for Medium and 0.28 for Hard tasks. These findings demonstrate the potential of LLMs to streamline the data science process, even for users without advanced coding expertise, with greater reliance on LLMs for easier tasks and more mixed results for more complex ones.
@@
The quantitative results from the user survey are summarized in Fig. 4e, where we grouped the questions into four main categories. The average user ratings for each category are: Output Quality (3.4), Support & Integration (3.0), System Complexity (3.5), and System Usability (4.0). These ratings suggest that, overall, users had a positive experience using the platform. Additionally, we collected qualitative feedback (Fig. 4f), where one user expressed a strong interest in continuing to use AI for research on their own data, highlighting the platform’s practical utility. Another user acknowledged the platform’s value in helping them learn programming and data analysis, underscoring its potential as an educational tool for those with limited coding experience. These insights reinforce the platform’s ability to enhance both productivity and learning in data science workflows.
3 Discussion
@@
In collaboration with medical experts, data scientists play a pivotal role in analyzing complex datasets, such as real-world patient data, to derive insights that improve patient care and inform evidence-based medicine. However, the rising demand for data science expertise, combined with the limited availability of skilled professionals, has created a bottleneck, slowing progress and hindering the full potential of data-driven clinical research. This shortage is restricting the ability to fully harness the vast amount of data available for advancing clinical research.
@@
Large language models (LLMs) have emerged as powerful generalist artificial intelligence (AI) capable of following human instructions to perform a wide range of tasks in medicine [37, 38, 39]. In parallel, LLMs have demonstrated strong capabilities in solving coding challenges [5], completing software engineering tasks [13], and performing basic data analysis [20]. These advancements suggest that LLMs hold great promise for streamlining data science projects in clinical research, a potential that has not yet been fully explored.
@@
The primary goal of this study is to thoroughly evaluate the performance of cutting-edge LLMs in handling complex clinical research data and performing data science programming tasks. To achieve this, we developed a comprehensive coding benchmark, CliniDSBench, comprising 39 published clinical studies and 293 diverse, practical, and high-quality data analysis tasks in both Python and R. Based on these benchmarks, we found that current LLMs are not yet capable of fully automating data science tasks. At their first attempts, LLMs successfully solved only 40%-80% of Easy tasks, 15%-40% of Medium tasks, and 5%-15% of Hard tasks. This highlights the necessity of human oversight and post-processing to prevent misinformation and incorrect results when relying on LLMs for clinical data analysis.
@@
Though imperfect, we found that much of the LLM-generated code was quite close to the correct solution. This observation motivated us to explore advanced adaptation methods to improve LLM performance further. On the one hand, we found that LLMs could self-correct a significant portion of erroneous code, leading to substantial improvements over their initial attempts. On the other hand, involving human experts more directly in the process, such as by providing concrete, step-by-step plans for data analysis tasks, resulting in the best performance across all adaptation strategies.
@@
Beyond automatic testing, we conducted a user study using our developed interface, which integrates LLMs into the data science workflow. The study revealed that users heavily relied on LLM-generated code when crafting their final solutions, validating the effectiveness of LLMs in streamlining the coding process. This workflow typically followed a pattern where LLMs provided an initial solution, users collaborated with the LLMs for debugging, and then the users refined the final solution. The platform was highly appreciated by users, not only for its practical utility in accelerating data science tasks but also for its educational value in helping them improve their programming and data analysis skills.
@@
This study has several limitations. First, to ensure the quality of the benchmark, we manually created all the questions and solutions for the analysis tasks, which restricted our ability to scale the benchmark to cover more studies. A larger dataset with more coding tasks would not only enhance evaluation but could also be used for training LLMs specifically for data science tasks. Second, the user study results may be biased, as the participants were primarily medical doctors who, while knowledgeable in their domain, had varying levels of coding proficiency. The usage patterns might differ significantly if data scientists were the users, as they possess more advanced coding skills but know less about medicine. Third, the patient-level data in our testing set are publicly available, but privacy risks must be carefully considered when deploying LLMs for real-world clinical data analysis. A recommended approach would be to separate the environment running code on sensitive patient data from the environment where LLMs are used. LLMs should only access the dataset schema or global statistics without access to individual patient data. Finally, the patient data used in our testing set were relatively clean, standardized, and semantically meaningful. In real-world scenarios, data can be messier and more varied, which could affect LLM performance. Future work should explore strategies to handle less structured, real-world data effectively.
@@
The findings from our study show that while LLMs are not yet capable of fully automating clinical research data science tasks, they can be valuable tools when used in collaboration with human experts. This human-AI partnership can lead to the creation of effective coding solutions, boost productivity, potentially accelerate drug development, and improve patient outcomes. However, testing this hypothesis requires future prospective studies in clinical research to confirm the practical impact of such collaborations.
4 Methods
4.1 Dataset curation
@@
We created the testing dataset, referred to as CliniDSBench, based on published clinical studies and their associated patient-level datasets from cBioPortal [23]. cBioPortal is a comprehensive database for cancer and genomics research, providing access to hundreds of studies with linked patient data. These datasets encompass various modalities, including clinical data, clinical sample data, mutations, copy number alterations, structural variants, RNA, mRNA, and tumor miRNA, among others. This setup ensures that the coding tasks in CliniDSBench are closely aligned with the real-world challenges faced in clinical research data science, using authentic data and analysis tasks.
@@
4.2 Large language models and adaptation methods
We investigated a diverse range of transformer-based large language models (LLMs) for data science code generation tasks, focusing on cutting-edge proprietary models. These include OpenAI’s GPT-4o [24] and GPT-4o-mini [25], Google’s Gemini-Pro and Gemini-Flash [28], as well as Anthropic’s Opus-3 [27] and Sonnet-3.5 [26]. Each of these models is a flagship proprietary LLM known for its strong performance in medical and biomedical tasks. Additionally, all these models feature long context windows, enabling them to handle large inputs efficiently: GPT-4o and GPT-4o-mini support up to 128K tokens, Gemini-Pro up to 2M tokens, Gemini-Flash up to 1M tokens, and both Opus-3 and Sonnet-3.5 support up to 200K tokens. This extended context capacity is essential for processing complex datasets and tasks typical in clinical research and data science.
No open-source code LLMs were included in this study for several key reasons. Firstly, most open-source code LLMs have limited context lengths, typically ranging from 2K to 8K tokens, which is insufficient for many of the data science tasks in our dataset. These tasks not only require input questions but also detailed dataset schema descriptions, sometimes spanning multiple tables with hundreds of columns. Secondly, previous studies have shown that open-source code LLMs significantly underperform compared to proprietary models, even on simpler tasks. For instance, in DS-1000 [20], proprietary models like Codex [41] outperformed open-source models such as CodeGen [42] and InCoder [43] by four to five times. Similarly, in BioCoder [22], GPT-4 achieved a Pass@1 rate of approximately 40%, while open-source models like StarCoder, even at 15.5 billion parameters [8], scored below 10%, despite fine-tuning. Given these findings, the proprietary LLMs used in our study can be considered to represent the upper bound of current LLM performance.
@@ 4.3 Experimental setup
We established a sandbox environment to enable the automatic execution and testing of code generated by LLMs. This was accomplished by creating a standardized Docker image that hosts both Python and R environments, allowing scripts in either language to be run via the command line. We utilized Pipenv to manage the Python environment, installing packages like pandas and matplotlib. Similarly, for R, we defined necessary packages such as dplyr and survival to be installed when building the image. The sandbox interface dynamically builds Docker containers based on the defined image, accepts code strings from LLMs, converts them into Python or R scripts, and then executes them. The sandbox also accepts dataset uploads, enabling parallel real-time code execution without impacting the main experimental environment. This setup ensures a controlled and isolated environment for running and testing LLM-generated code safely and efficiently.
@@
For the user study, we developed a platform to facilitate human-AI collaborative coding for data science projects, as illustrated in Extended Fig. 1. The platform is designed to relieve users from setting up their coding environment and provide code suggestions based on natural language requests, while enabling real-time code execution and feedback. The primary window features a user input box where users can choose from various platform commands, which are categorized into two types: brainstorming and programming. In the brainstorming mode, users can interact with the LLM assistant to search medical publications from PubMed or perform general searches via Google. They can also collaborate with LLMs to develop plans for data analysis tasks. In the programming mode, users can either ask the LLM to generate code from scratch or request improvements or corrections to existing code. Users have the option to generate code in either R or Python. Once the code is generated, users can execute it within a sandbox environment. The platform provides execution logs and any produced artifacts, such as figures, directly to the frontend, allowing users to receive immediate feedback on the results. This process maximizes the utility of LLMs by enabling users to collaboratively plan data analyses and then guide LLMs to generate accurate code solutions. In the second window, users can select patient datasets and apply their generated data analysis code to gain insights. The platform allows users to preview tables, columns, and values from the selected dataset, providing a streamlined experience for conducting data science tasks in a collaborative, AI-assisted environment.
@@
4.4 Evaluation metrics and statistical analysis
@@
Figure 1: Overview of the developed clinical research data science platform. a, code generation panel where users provide their requests, read the generated code, and switch to execution results. Users can view the previously sent requests and switch back if desired. b, Users can select one from a list of datasets to analyze. For each dataset, users can preview the content.
Figure 2: An example of Python coding task with the input question, prefix code, and testing cases.
Figure 3: Prompt for the Vanilla method in code generation.
Figure 4: Prompt for the Manual method in code generation.
Figure 5: Prompt for the AutoPrompt method in code generation.
Figure 6: Prompt for the Self-reflection method in code debugging and improvement.
Figure 7: Prompt for extracting and computing the number of semantic lines for estimating coding task difficulty.

### v2

@@ Results
The interface supports real-time interactions, allowing users to generate and execute code in a sandbox environment with instant visualizations. This removes the need for users to handle complex prompt crafting or manually switch between chat sessions and coding platforms like Jupyter Notebook. By simplifying the data science workflow, the platform empowers users with minimal coding expertise to perform complex data science tasks.
@@
In our user study, we involved five medical researchers with varying levels of coding expertise. Each participant was assigned three studies [33, 34, 35], with approximately 10 coding tasks per study (Fig. 5d). Users worked with LLMs on our platform to complete these tasks and submitted their solutions once their code passed all the test cases. The difficulty levels of the tasks were quantified, with the distribution shown in Fig. 5b. During the study, we tracked two core actions: code generation and code improvement (debugging) requests. The statistics of these user behaviors are depicted in Fig. 5b, where most users completed the first two studies, and a few tackled the third. After the study, we analyzed the logs to compare the LLM-generated code with the final code solutions submitted by the users (Fig. 5a). Additionally, we conducted a survey to gather their feedback on the platform and their experience working with LLMs. The survey questions were built based on the Health Information Technology Usability Evaluation Scale (Health-ITUES) [36].
@@
The results of the code comparison analysis are presented in Fig. 5c, showing the distribution of the proportion of user-submitted code derived from LLM-generated solutions. We found that a significant portion of the user-submitted code was drawn from LLM-generated code. For Easy tasks, the median proportions were 0.88, 0.87, and 0.84 across the three studies, indicating that users heavily relied on LLM-provided solutions when crafting their final submissions. For Medium and Hard tasks, the ratios were generally lower: in Study 1, the proportions were 0.44 for Medium tasks and 0.96 for Hard tasks, while in Study 2, the proportions were 0.75 for Medium and 0.28 for Hard tasks. These findings demonstrate the potential of LLMs to streamline the data science process, even for users without advanced coding expertise, with greater reliance on LLMs for easier tasks and more mixed results for more complex ones.
@@
The quantitative results from the user survey are summarized in Fig. 5e, where we grouped the questions into four main categories. The average user ratings for each category are: Output Quality (3.4/5), Support & Integration (3.0/5), System Complexity (3.5/5), and System Usability (4.0/5). These ratings suggest that, overall, users had a positive experience using the platform. Additionally, we collected qualitative feedback (Fig. 5f), where one user expressed a strong interest in continuing to use AI for research on their own data, highlighting the platform’s practical utility. Another user acknowledged the platform’s value in helping them learn programming and data analysis, underscoring its potential as an educational tool for those with limited coding experience. These insights reinforce the platform’s ability to enhance both productivity and learning in data science workflows.
Discussion
@@
In collaboration with medical experts, data scientists play a pivotal role in analyzing complex datasets, such as real-world patient data, to derive insights that improve patient care and inform evidence-based medicine. However, the rising demand for data science expertise, combined with the limited availability of skilled professionals, has created a bottleneck, slowing progress and hindering the full potential of data-driven biomedical research. This shortage is restricting the ability to fully harness the vast amount of data available for advancing biomedical research.
@@
Large language models (LLMs) have emerged as powerful generalist AI capable of following human instructions to perform a wide range of tasks in medicine [37, 38, 39]. In parallel, LLMs have demonstrated strong capabilities in solving coding challenges [5], completing software engineering tasks [13], and performing basic data analysis [20]. These advancements suggest that LLMs hold great promise for streamlining data science projects in biomedical research, a potential that has not yet been fully explored.
@@
The primary goal of this study is to thoroughly evaluate the performance of cutting-edge LLMs in handling complex biomedical research data and performing data science programming tasks. To achieve this, we developed a comprehensive coding benchmark, BioDSBench, comprising 39 published medical and biomedical studies and 293 diverse, practical, and high-quality data analysis tasks in both Python and R. Based on these benchmarks, we found that current LLMs are not yet capable of fully automating data science tasks. At their first attempts, LLMs successfully solved only 40-80% of Easy tasks, 15-40% of Medium tasks, and 5-15% of Hard tasks. This highlights the necessity of human oversight and post-processing to prevent mistakes and incorrect results when relying on LLMs for biomedical data analysis.
@@
Though imperfect, we found that much of the LLM-generated code was quite close to the correct solution. This observation motivated us to explore advanced adaptation methods to improve LLM performance further. On the one hand, we found that LLMs could self-correct a significant portion of erroneous code, leading to substantial improvements over their initial attempts. In particular, involving human experts more directly in the process, such as by providing concrete, step-by-step plans for data analysis tasks, resulting in the best performance across all adaptation strategies.
@@
Beyond automatic testing, we conducted a user study using our developed interface, which integrates LLMs into the data science workflow. The study revealed that users heavily relied on LLM-generated code when crafting their final solutions, validating the effectiveness of LLMs in streamlining the coding process. This workflow typically followed a pattern where LLMs provided an initial solution, users collaborated with the LLMs for debugging, and then the users refined the final solution. Our platform was appreciated by users, not only for its practical utility in accelerating data science tasks but also for its educational value in helping them improve their programming and data analysis skills.
@@
This study has several limitations. First, to ensure the quality of the benchmark, we manually created all the questions and solutions for the analysis tasks, which restricted our ability to scale the benchmark to cover more studies. A larger dataset with more coding tasks would not only enhance evaluation but could also be used for training LLMs specifically for data science tasks. Second, the user study results may be biased, as the participants were primarily medical researchers who, while knowledgeable in their domain, had varying levels of coding proficiency. The usage patterns might differ significantly if data scientists were the users, as they possess more advanced coding skills but know less about medicine. Third, the patient-level data in our testing set are publicly available, but privacy risks must be carefully considered when deploying LLMs for real-world biomedical data analysis. A recommended approach would be to separate the environment running code on sensitive patient data from the environment where LLMs are used. LLMs should only access the dataset schema or global statistics without access to individual patient data. Finally, the patient data used in our testing set were relatively clean, standardized, and semantically meaningful. In real-world scenarios, data can be messier and more varied, which could affect LLM performance. Future work should explore strategies to handle less structured, real-world data effectively.
@@
The findings from our study show that while LLMs are not yet capable of fully automating biomedical data science tasks, they can be valuable tools when used in collaboration with human experts. This human-AI partnership can lead to the creation of effective coding solutions, boost productivity, and accelerate biomedical research.
Methods
Dataset curation
@@ Methods
We created the testing dataset, referred to as BioDSBench, based on published medical studies and their associated patient-level datasets from cBioPortal [23]. cBioPortal is a comprehensive database for cancer and genomics research, providing access to hundreds of studies with linked patient data. These datasets encompass various modalities, including clinical data, clinical sample data, mutations, copy number alterations, structural variants, RNA, mRNA, and tumor miRNA, among others. This setup ensures that the coding tasks in BioDSBench are closely aligned with the real-world challenges faced in biomedical data science, using authentic data and analysis tasks.
@@
Large language models and adaptation methods
We investigated a diverse range of large language models (LLMs) for data science code generation tasks, focusing on cutting-edge proprietary models. These include OpenAI’s GPT-4o [24] and GPT-4o-mini [25], Google’s Gemini-Pro and Gemini-Flash [28], as well as Anthropic’s Opus-3 [27] and Sonnet-3.5 [26]. Each of these models is a flagship proprietary LLM known for its strong performance in medical and biomedical tasks. Additionally, all these models feature long context windows, enabling them to handle large inputs efficiently: GPT-4o and GPT-4o-mini support up to 128K tokens, Gemini-Pro up to 2M tokens, Gemini-Flash up to 1M tokens, and both Opus-3 and Sonnet-3.5 support up to 200K tokens. This extended context capacity is essential for processing complex datasets and tasks typical in biomedical data science.
No open-source code LLMs were included in this study for several reasons. First, most open-source code LLMs have limited context lengths, typically ranging from 2K to 8K tokens, which is insufficient for many of the data science tasks in our dataset. These tasks not only require input questions but also detailed dataset schema descriptions, sometimes spanning multiple tables with hundreds of columns. Second, previous studies have shown that open-source code LLMs significantly underperform compared to proprietary models, even on simpler tasks. For instance, in DS-1000 [20], proprietary models like Codex [41] outperformed open-source models such as CodeGen [42] and InCoder [43] by four to five times. Similarly, in BioCoder [22], GPT-4 achieved a Pass@1 rate of approximately 40%, while open-source models like StarCoder, even at 15.5 billion parameters [8], scored below 10%, despite fine-tuning. Given these findings, the proprietary LLMs used in our study can be considered to represent the upper bound of current LLM performance.
@@
We created a sandbox environment to enable the automatic execution and testing of code generated by LLMs. This was accomplished by creating a standardized Docker image that hosts both Python and R environments, allowing scripts in either language to be run via the command line. We utilized Pipenv to manage the Python environment, installing packages like pandas and matplotlib. Similarly, for R, we defined necessary packages such as dplyr and survival to be installed when building the image. The sandbox interface dynamically builds Docker containers based on the defined image, accepts code strings from LLMs, converts them into Python or R scripts, and then executes them. The sandbox also accepts dataset uploads, enabling parallel real-time code execution without impacting the main experimental environment. This setup ensures a controlled and isolated environment for running and testing LLM-generated code safely and efficiently.
@@
For the user study, we developed a platform to facilitate human-AI collaborative coding for data science tasks, as illustrated in Fig. 4. The platform is designed to relieve users from setting up their coding environment and provide code suggestions based on natural language requests, while enabling real-time code execution and feedback. The primary window features a user input box where users can choose from various platform commands, which are categorized into two types: brainstorming and programming. In the brainstorming mode, users can interact with the LLM assistant to search medical publications from PubMed or perform general searches via Google. They can also collaborate with LLMs to develop plans for data analysis tasks. In the programming mode, users can either ask the LLM to generate code from scratch or request improvements or corrections to existing code. Users have the option to generate code in either R or Python. Once the code is generated, users can execute it within a sandbox environment. The platform provides execution logs and any produced artifacts, such as figures, directly to the frontend, allowing users to receive immediate feedback on the results. This process maximizes the utility of LLMs by enabling users to collaboratively plan data analyses and then guide LLMs to generate accurate code solutions. In the second window, users can select patient datasets and apply their generated data analysis code to gain insights. The platform allows users to preview tables, columns, and values from the selected dataset, providing a streamlined experience for conducting data science tasks in a collaborative, AI-assisted environment.
@@
Evaluation metrics and statistical analysis
@@
Data availability
The curated data science tasks with the reference answers and testing cases in BioDSBench can be accessed via https://huggingface.co/datasets/zifeng-ai/BioDSBench. The anonymized patient data where these data analyses are performed are available from the cBioPortal website (https://www.cbioportal.org/datasets) and the UCSC Xena website (https://xenabrowser.net/datapages/).
Code availability
Code for implementing and experimenting with the proposed methodology is available at https://github.com/RyanWangZf/BioDSBench. The human-AI collaborative biomedical data science programming platform can be accessed via a web-based app (https://www.trialmindapis.com/api/data-science) and can be accessed per request via https://keiji.ai/contact.html. The demonstration video can be accessed via https://www.youtube.com/watch?v=c5ZJsFXQ_B0.
Figure 2: Assessment of different models and adaptation methods in automating biomedical data science tasks. a, the inputs for LLMs to generate the code and the associated evaluation process. b, the pass@5 of three LLMs with varying temperatures across difficulty levels in the Python coding dataset. c, the proportions of the reference solution code that can be drawn directly from the LLM-generated code. d and e show the pass@1 of six LLMs across difficulty levels in Python and R coding datasets, respectively.
Figure 4: Overview of the developed biomedical data science platform. a, code generation panel where users provide their requests, read the generated code, and switch to execution results. Users can view the previously sent requests and switch back if desired. b, Users can select one from a list of datasets to analyze. For each dataset, users can preview the content.